# Face Mask Removal with Region-attentive Face Inpainting

### v1

@@ 4.3 Performance Evaluation
We test all the methods on the validation data ($19,026$ images in total) of the Masked-Faces dataset. For quantitative comparison, we employ the metrics of SSIM [48], PSNR and $\ell_{1}$ loss. SSIM [48] is used for measuring the quality of digital images and videos. It consists of three components: luminance ($l$), contrast ($c$) and structure ($s$). Assume $x_{1}$ and $x_{2}$ are two non-negative image signals. The SSIM between $x_{1}$ and $x_{2}$ can be computed as follows:
$l(x_{1},x_{2})=(2\mu_{x1}\mu_{x2}+c_{1})/(\mu_{x1}^{2}+\mu_{x2}^{2}+c_{1})$ (11)
$c(x_{1},x_{2})=(2\sigma_{x1}\sigma_{x2}+c_{2})/(\sigma_{x1}^{2}+\sigma_{x2}^{2}+c_{2})$ (12)
$s(x_{1},x_{2})=(\sigma_{x_{1}x_{2}}+c_{3})/(\sigma_{x1}\sigma_{x2}+c_{3})$ (13)
$SSIM(x_{1},x_{2})=l(x_{1},x_{2})\cdot c(x_{1},x_{2})\cdot s(x_{1},x_{2})$ (14)
where $\mu_{x1}$ and $\mu_{x2}$ are the means of $x_{1}$ and $x_{2}$, respectively, $\sigma_{x1}$ and $\sigma_{x2}$ are the variance of $x_{1}$ and $x_{2}$, respectively, and $c_{1}$, $c_{2}$ and $c_{3}$ are variables for stabilizing the division.
We calculate the SSIM [48] between the recovered image and the ground truth. The results are provided in Tab. 1. As can be seen, our proposed method outperforms all the other works in terms of the SSIM [48], PSNR and $\ell_{1}$ loss.

### v2

@@ 4.3 Performance Evaluation
We test all the methods on the validation data ($19,026$ images in total) of the Masked-Faces dataset. For quantitative comparison, we employ the metrics of SSIM [49], PSNR and $\ell_{1}$ loss. We calculate the SSIM [49] between the recovered image and the ground truth. The results are provided in Tab. 1. As can be seen, our proposed method outperforms all the other works in terms of the SSIM [49], PSNR and $\ell_{1}$ loss.
@@
4.4 Face Recognition Results
For this experiment, we employ two different face recognition networks, namely MobileFaceNet [4] and ArcFace [6], to show (i) the degradation of face recognition performance when masked face images are used, and (ii) the improvement after face inpainting. Table 2 reveals two key observations. First, face masks deteriorate the face recognition performance considerably (by 7.85% on average for 2 networks), whereas using the inpainted images, from all methods, results in higher face recognition accuracy, with images inpainted by our approach resulting in the highest accuracy. Second, the accuracy comparison across all inpainting approaches show that our approach has a better identity-preserving ability to facilitate face recognition, which can be explained by our method’s ability to recover the masked region with better-quality details and higher similarity to ground truth.
Input MobileFaceNet [4] ArcFace [6]
Grnd Truth Img. 77.61 76.44
Masked Image 68.31 70.04
GatedConv [54] 71.72 72.08
CSA [31] 72.13 72.54
RFR [26] 71.34 73.71
RN [55] 72.93 72.37
Ours 75.67 73.87
Table 2: Face Recognition Accuracy (%). Among the inpainted images, the ones obtained with our method result in the highest face recognition accuracy.